# Omni-Effects: Unified and Spatially Controllable Visual Effects Generation

### v3

@@ Appendix C Implementation Details
C.1 Training Data Augmentation
@@
C.2 Non-Uniform Timestep Sampling
@@
C.3 Dual-phase Training Strategy
@@ Appendix D Metrics
With the development of diffusion-based generative models Chu et al. [2025], Lan et al. [2025], Xu et al. [2025a], a series of evaluation benchmarks have emerged to assess the quality of generated content Ling et al. [2025], Feng et al. [2025], Chen et al. [2025], Huang et al. [2024a]. However, existing benchmarks struggle to meet the evaluation requirements for controllable VFX.
Therefore, we propose an evaluation framework comprising three components: VFX detection, VFX controllable estimation, and video motion estimation, with the specific architecture illustrated in the Figure 14. These components correspond to the calculation processes of the metrics Effect Occurrence Rate (EOR), Effect Controllability Rate (ECR), and Regional Dynamic Degree (RDD).
D.1 Visual Effects Detection
To determine whether the desired visual effect is present in the generated videos, we leverage the Gemini 2.5 Comanici et al. [2025] large multimodal model as an evaluation assistant. Specifically, for each test video, we pair it with a predefined prompt template that explicitly describes the target effect, and input both into Gemini. We then request Gemini to provide a binary answer as to whether the specified effect has appeared in the video. Each query is repeated three times, and the most frequently occurring answer among the three is selected as the final result.
@@ D.1 Visual Effects Detection
D.2 Visual Effects Controllability Estimation
@@
D.3 Video Motion Detection
The Regional Dynamic Degree (RDD) measures the strength of motion caused by VFX within mask-specified region of a video. We use the RAFT algorithm Teed and Deng [2020] to estimate optical flow between consecutive frames. Given a binary mask for the region of interest, we calculate the mean motion magnitude within this region as follows:
@@ D.3 Video Motion Detection
Appendix E Experiments Results Details
E.1 Qualitative Results of LoRA MoE
E.2 Qualitative Results of single-VFX control
E.3 User Study of Multi-VFX Generation
Appendix F More Ablation Study
F.1 VFX-combination training

### v4

@@ Appendix C Implementation Details
Training Data Augmentation
@@
Non-Uniform Timestep Sampling
@@
Dual-phase Training Strategy
@@ Appendix D Metrics
Our evaluation framework comprises three components: VFX detection, VFX controllable estimation, and video motion estimation, with the specific architecture illustrated in the Figure 14. These components correspond to the calculation processes of the metrics Effect Occurrence Rate (EOR), Effect Controllability Rate (ECR), and Regional Dynamic Degree (RDD).
Visual Effects Detection
To determine whether the desired visual effect is present in the generated videos, we leverage the Gemini 2.5 (comanici2025gemini) large multimodal model as an evaluation assistant. Specifically, for each test video, we pair it with a predefined prompt template that explicitly describes the target effect, and input both into Gemini. We then request Gemini to provide a binary answer as to whether the specified effect has appeared in the video. Each query is repeated three times, and the most frequently occurring answer among the three is selected as the final result.
@@
Visual Effects Controllability Estimation
@@
Video Motion Detection
The Regional Dynamic Degree (RDD) measures the strength of motion caused by VFX within mask-specified region of a video. We use the RAFT algorithm (teed2020raft) to estimate optical flow between consecutive frames. Given a binary mask for the region of interest, we calculate the mean motion magnitude within this region as follows:
@@
Appendix E Experiments Results Details
Qualitative Results of LoRA MoE
Qualitative Results of single-VFX control
User Study of Multi-VFX Generation
Appendix F More Ablation Study
VFX-combination training
@@ Appendix F More Ablation Study
Appendix G Limitation
As the number of combined VFX categories increases (especially when ), the performance tends to decrease. We attribute this phenomenon to the lack of real multi-VFX data, which leads to a bias between training and inference. In future work, we plan to collect more high-quality multi-VFX data to improve the stability of our method.